OpenStudy (anonymous):

find a set of parametric equations for the line or conic. Ellipse: Vertices: (+/-5,0) & Foci: (+/-4,0)

OpenStudy (anonymous):

do you know what the center is?

OpenStudy (anonymous):



OpenStudy (anonymous):

no it provides us with that

OpenStudy (anonymous):

the center of the ellipse is the two foci isn't it?

OpenStudy (anonymous):

right i know, my question really meant "given the information, what is the center?"

OpenStudy (anonymous):

no there is one center for an ellipse two foci, but one center it is half way between the foci, so in this case it is right at the origin \((0,0)\)

OpenStudy (anonymous):

the general form of an ellipse is \[\frac{(x-h)^2}{a^2}+\frac{(y-h)^2}{b^2}=1\] but in your example, since the center is \((0,0)\) it is \[\frac{x^2}{a^2}+\frac{y^2}{b^2}=1\] and your job now is to find \(a\) and \(b\)

OpenStudy (anonymous):

there asking for two parametric equations and would would put them into these x-h+acos theta y=k+bcos theta i just don't know what to plug in where

OpenStudy (anonymous):

oh i see you can ignore the \(h\)and \(k\) part because the center is \((0,0)\) and so \(h=k=0\)

OpenStudy (anonymous):

so it is just going to be \[x=a\cos(\theta)\] and \[y=b\sin(\theta)\]

OpenStudy (anonymous):

since the ellipse travels through the point \((5,0)\) you know it is \[x=5\cos(\theta)\]

OpenStudy (anonymous):

does the negative have any influence?

OpenStudy (anonymous):

that way when \(\theta=0\) you get \(5\cos(0)=5\) and \(5\sin(0)=0\)

OpenStudy (anonymous):

no not really \(\theta\) will go form \(0\) to \(2\pi\) so you will get negative values that way the only work you need to do here is to find where the ellipse hits the \(y\) axis so you can put that number for \(b\) in \(b\sin(\theta)\)

OpenStudy (anonymous):

and you do that by noting that \(a^2=b^2+c^2\) or in your case \[5^2=b^2+4^2\] making \[b=3\]

OpenStudy (anonymous):

your parametric equations are therefore \[x=5\cos(\theta), y=3\sin(\theta)\]

OpenStudy (anonymous):

and the 4 means basically nothing?

OpenStudy (anonymous):

oh no it means something it tells you what the foci are but in this case it tells you how to find \(b\)

OpenStudy (anonymous):

ok. thanks!!!!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

how would you solve this type of problem but for a hyperbola?